# Pointwise convergence of ergodic averages with Möbius weight

### v1

@@ 3.1. Asymptotic notation, indicators and averages
We also use the double-logarithmic averages $\mathbb{E}^{\log\log}_{n\leq N}$, defined in (1.3).
For a real number $N\geq 1$, we denote $[N]\coloneqq\{n\in\mathbb{N}\colon\,\,n\leq N\}$. For integers $m,n$, we denote their greatest common divisor by $(m,n)$ and write $m\mid n^{\infty}$ to mean that $m\mid n^{k}$ for some natural number $k$. Unless otherwise specified, all our sums and averages run over the integers, with the exception that the variable $p$ is reserved for primes.
@@ 3.2. Gowers norms
As is well known (see for example [12, Appendix B]), for $s\geq 2$ the $U^{s}[N]$ norm is indeed a norm and for $s=1$ it is a seminorm, and the function $s\mapsto\|f\|_{U^{s}[N]}$ is increasing.
@@
For the $\widetilde{U}^{s}(\mathbb{Z})$ norms we have the Gowers–Cauchy–Schwarz inequality (see for example [26, (4.2)]), which states that, for any functions $(f_{\omega})_{\omega\in\{0,1\}^{s}}$ from $\mathbb{Z}$ to $\mathbb{C}$ with finite support, we have
@@
We remark that, as is well known, for $s=2$ the $u^{s}[N]$ and $U^{s}[N]$ norms are equivalent (up to polynomial losses), but for all $s\geq 3$ the $u^{s}[N]$ norm is weaker in the sense that the $U^{s}[N]$ norm controls the $u^{s}[N]$ norm but not vice versa; see [14][11, Section 4].
@@ 3.3. Van der Corput’s inequality
where $\lfloor x\rfloor$ is the floor of $x$ (This should not be confused with the Möbius function $\mu$). For an integer $h$, we define the differencing operator $\Delta_{h}$ by setting $\Delta_{h}f(x)=\overline{f(x+h)}f(x)$ for any $f\colon\mathbb{Z}\to\mathbb{C}$ and $x\in\mathbb{Z}$.
@@ 3.4. Vinogradov’s Fourier expansion
This follows from [27, Lemma 12] (taking $r=1$ there). ∎
@@ 4. Polynomial generalised von Neumann theorems
where $P_{1},\ldots,P_{k}\in\mathbb{Z}[y]$ and $\theta\colon[N]\to\mathbb{C}$ is a weight function (which we take in applications to be the Möbius function) and $f_{0},f_{1},\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ are functions supported on $[-M,M]$ (with $M\asymp N^{\max_{j}(\deg P_{j})}$). Thus, we bound the expression (4.1) in terms of some Gowers norm (or related norm) of $\theta$ or $f_{0}$. It is important for the proofs of our main theorems that the obtained results are quantitative, with polynomial dependencies. The two main results of this section (which are of independent interest) are Theorems 4.1 and 4.2.
The first main result of this section states that if $P_{1},\ldots,P_{k}$ have distinct degrees, then the polynomial counting operator (4.1) is bounded in terms of the $u^{s}[N]$ norm of the weight $\theta$ for some $s$, with polynomial dependencies. It is important to have the $u^{s}[N]$ norm rather than the $U^{s}[N]$ norm here, since for the former we have much better quantitative bounds in the case of the Möbius function.
@@ 4.1. Transferring to locally linear functions
Proposition 4.3 may be compared with, and is motivated by, the work of Peluse and Prendiville [24, Theorem 1.5], where in the case $k=2$, $\theta\equiv 1$ and $P_{1}(y)=y,P_{2}(y)=y^{2}$, it is proved that $f_{0},f_{1},f_{2}$ can be replaced more strongly with major arc locally linear phase functions. In the more general setup of Proposition 4.3, it is not possible to reduce to major arc locally linear phase functions.
@@
This will follow from [23, Theorem 3.3] after some reductions. It suffices to show that for each $j\in\{0,1\}$ there exists $11\leq B\ll_{d}$ 1 such that (4.3) holds, since we may increase $B$ if necessary.
@@
Now, applying [^2] In  [23, Theorem 3.3], the functions $f_{i}$ are assumed to be supported on $[CN^{d}]$ instead of $[-CN^{d},CN^{d}]$, but this makes no difference in the argument. [23, Theorem 3.3] (with $\eta$ in place of $\delta$) we see that there exists some $11\leq B\ll_{d}$ 1 such that
@@
Now, extend the definition of $\alpha_{m}$ to all of $\mathbb{Z}$ by letting $\alpha_{m}=\alpha_{m^{\prime}}$, where $m^{\prime}$ is the largest element of $N^{\prime}\mathbb{Z}$ that is at most $m$. Then, define the locally linear phase function $\phi_{0}(m)=e(\alpha_{m}m)$, which has resolution $N^{\prime}$. We now have
@@
$G(m)\coloneqq\mathbb{E}_{n\in[N]}\theta(n)\phi_{0}(m-P_{1}(n))f_{2}(m+P_{2}(n)-P_{1}(n))\cdots f_{k}(m+P_{k}(n)-P_{1}(n))$
@@ 4.2. A circle method bound
We are now ready to prove the claimed estimate for (4.1) for distinct degree polynomials.
@@
For the proof of Theorem 4.2, we need the following generalised von Neumann theorem for arithmetic progressions. This is well known (see for example [9, Lemma 2]), although the result is typically presented for functions defined on a cyclic group.
@@
(4.22) $\begin{split}|S|^{2}&\ll\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{Q\in\mathcal{Q}}\Delta_{h}f_{Q}(m+Q(n))\\ &=\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{\widetilde{Q}\in\mathcal{Q}_{h}}\widetilde{f}_{\widetilde{Q}}(m+\widetilde{Q}(n)),\end{split}$
where each $\widetilde{f}_{\widetilde{Q}}$ is equal to $f_{Q}$ or $\overline{f_{Q}}$ for some $Q\in\mathcal{Q}$ and $\mathcal{Q}_{h}$ is the collection (of size $\leq 2|\mathcal{Q}|$ and degree $\leq\deg\mathcal{Q}$) given by
@@
where $\widetilde{f}_{\widetilde{Q},a}=\widetilde{f}_{\widetilde{Q}}(\cdot+a)$. Note that by the mean value theorem for all $\widetilde{Q}\in\mathcal{Q}_{h}$ we have $\max_{n\in[N]}|\widetilde{Q}(n)|\ll_{\deg\mathcal{Q}}N^{d_{h}}$, so for all $n\in[N]$ we have $\widetilde{Q}(n)\in[-C^{\prime}N^{d_{h}},C^{\prime}N^{d_{h}}]$ for some $C^{\prime}\ll_{\deg\mathcal{Q}}C$. Now, by the induction assumption (and the fact that the functions $\widetilde{f}_{Q,jN^{d}_{h}}$ may be assumed to be supported on $[-(C^{\prime}+1)N^{d_{h}},(C^{\prime}+1)N^{d_{h}}]$), we conclude that for some natural number $s^{\prime}\ll_{|\mathcal{Q}|,\deg{Q}}$ we have
@@
5. Lemmas for the main proofs
5.1. Quantitative uniformity bounds for the Möbius function
In this subsection, we state the quantitative bounds for the uniformity of the Möbius function that we need for the proofs of the main theorems.
Lemma 5.1 (Quantitative $U^{k}$- and $u^{k}$-uniformity of $\mu$).
Let $k\in\mathbb{N}$, $A>0$ and $N\geq 3$.
  1. (i)
    We have
    $\|\mu\|_{U^{4}[N]}\ll_{A}(\log N)^{-A}.$
  2. (ii)
    We have
    $\|\mu\|_{U^{k}[N]}\ll_{k}(\log\log N)^{-c_{k}}$
    for some $c_{k}>0$.
  3. (iii)
    We have
    $\|\mu\|_{u^{k}[N]}\ll_{A,k}(\log N)^{-A}.$
@@ 5.1. Quantitative uniformity bounds for the Möbius function
Part (i) follows by combining recent work of Leng [22, Theorems 4 and 5] with the triangle inequality for the Gowers norms. Part (ii) was proven in [26, Theorem 1.4]. Recalling Definition 3.5 of the $u^{k}[N]$ norm, part (iii) follows from the Green–Tao Möbius and nilsequences estimate [13, Theorem 1.1] (but it would also be possible to prove it by more classical means by using estimates for bilinear sums of polynomial exponential phases). ∎
@@
Prior to [22], Leng proved in [21] that $\|\mu\|_{U^{3}[N]}\ll_{A}(\log N)^{-A}$. This result would already suffice for proving Theorem 1.2(i) for $k=2$; see Theorem  6.1.

### v2

@@ 3.1. Asymptotic notation, indicators and averages
For a real number $N\geq 1$, we denote $[N]\coloneqq\{n\in\mathbb{N}\colon\,\,n\leq N\}$. For integers $m,n$, we denote their greatest common divisor by $(m,n)$ and write $m\mid n^{\infty}$ to mean that $m\mid n^{k}$ for some natural number $k$. Unless otherwise specified, all our sums and averages run over the positive integers, with the exception that the symbol $p$ is reserved for primes.
@@ 3.2. Gowers norms
As is well known (see for example [16, Appendix B]), for $s\geq 2$ the $U^{s}[N]$ norm is indeed a norm and for $s=1$ it is a seminorm, and the function $s\mapsto\|f\|_{U^{s}[N]}$ is increasing.
@@
For the $\widetilde{U}^{s}(\mathbb{Z})$ norms we have the Gowers–Cauchy–Schwarz inequality (see for example [32, (4.2)]), which states that, for any functions $(f_{\omega})_{\omega\in\{0,1\}^{s}}$ from $\mathbb{Z}$ to $\mathbb{C}$ with finite support, we have
@@
We remark that, as is well known, for $s=2$ the $u^{s}[N]$ and $U^{s}[N]$ norms are equivalent (up to polynomial losses), but for all $s\geq 3$ the $u^{s}[N]$ norm is weaker in the sense that the $U^{s}[N]$ norm controls the $u^{s}[N]$ norm but not vice versa; see [18][14, Section 4].
@@ 3.3. Van der Corput’s inequality
where $\lfloor x\rfloor$ is the floor function of $x$ (This weight should not be confused with the Möbius function $\mu$). For an integer $h$, we define the differencing operator $\Delta_{h}$ by setting $\Delta_{h}f(x)=\overline{f(x+h)}f(x)$ for any $f\colon\mathbb{Z}\to\mathbb{C}$ and $x\in\mathbb{Z}$.
@@ 3.4. Vinogradov’s Fourier expansion
This follows from [33, Lemma 12] (taking $r=1$ there). ∎
@@ 4. Polynomial generalised von Neumann theorems
where $P_{1},\ldots,P_{k}\in\mathbb{Z}[y]$ and $\theta\colon[N]\to\mathbb{C}$ is a weight function (which in applications we take to be a multiplicative function) and $f_{0},f_{1},\ldots,f_{k}\colon\mathbb{Z}\to\mathbb{C}$ are functions supported on $[-M,M]$ (with $M\asymp N^{\max_{j}(\deg P_{j})}$). Thus, we bound the expression (4.1) in terms of some Gowers norm (or related norm) of $\theta$ or $f_{0}$. It is important for the proofs of our main theorems that the obtained results are quantitative, with polynomial dependencies. The two main results of this section (of independent interest) are Theorems 4.1 and 4.2; they are used for proving Theorems 1.6 and 1.2, respectively.
The first main result of this section states that if $P_{1},\ldots,P_{k}$ have distinct degrees, then the polynomial counting operator (4.1) is bounded in terms of the $u^{s}[N]$ norm of the weight $\theta$ for some $s$, with polynomial dependencies. It is important to have the $u^{s}[N]$ norm rather than the $U^{s}[N]$ norm here, since for the $U^{s}[N]$ norm we do not currently have a polynomial inverse theorem, meaning that the Siegel–Walfisz assumption from Definition 1.5 would be insufficient if we only had a bound in terms of these norms.
@@ 4.1. Transferring to locally linear functions
Proposition 4.3 may be compared with, and is motivated by, the work of Peluse and Prendiville [29, Theorem 1.5], where in the case $k=2$, $\theta\equiv 1$ and $P_{1}(y)=y,P_{2}(y)=y^{2}$, it is proven that $f_{0},f_{1},f_{2}$ can be replaced more strongly with major arc locally linear phase functions. In the more general setup of Proposition 4.3, it is not possible to reduce to major arc locally linear phase functions.
@@
This will follow from [28, Theorem 3.3] after some reductions. It suffices to show that for each $j\in\{0,1\}$ there exists $11\leq B\ll_{d}$ 1 such that (4.3) holds, since we may increase $B$ if necessary.
@@
Now, applying [^4] In  [28, Theorem 3.3], the functions $f_{i}$ are assumed to be supported on $[CN^{d}]$ instead of $[-CN^{d},CN^{d}]$, but this makes no difference in the argument. [28, Theorem 3.3] (with $\eta$ in place of $\delta$) we see that there exists some $11\leq B\ll_{d}$ 1 such that
@@
Now, extend the definition of $\alpha_{m}$ from $N^{\prime}\mathbb{Z}$ to all of $\mathbb{Z}$ by letting $\alpha_{m}=\alpha_{m^{\prime}}$, where $m^{\prime}$ is the largest element of $N^{\prime}\mathbb{Z}$ that is at most $m$. Then, define the locally linear phase function $\phi_{0}(m)=e(\alpha_{m}m)$, which has resolution $N^{\prime}$. We now have
@@
$G(m)\coloneqq\mathbb{E}_{n\in[N]}\theta(n)\phi_{0}(m-P_{1}(n))f_{2}(m+P_{2}(n)-P_{1}(n))\cdots f_{k}(m+P_{k}(n)-P_{1}(n)).$
@@ 4.2. A circle method bound
We are now ready to prove the claimed estimate for the operator (4.1) in the case of distinct degree polynomials.
@@
For the proof of Theorem 4.2, we need the following generalised von Neumann theorem for arithmetic progressions. This is well known (see for example [11, Lemma 2]), although the result is typically presented for functions defined on a cyclic group.
@@
(4.22) $\begin{split}|S|^{2}&\ll\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{Q\in\mathcal{Q}}\Delta_{h}f_{Q}(m+Q(n))\\ &=\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{Q\in\mathcal{Q}}f_{Q}(m+Q(n+h)-Q_{1}(n))\overline{f_{Q}}(m+Q(n)-Q_{1}(n))\\ &=\frac{1}{N^{d+1}}\sum_{h\in\mathbb{Z}}\mu_{N}(h)\sum_{m\in\mathbb{Z}}\sum_{n\in[N]}\Delta_{h}\theta(n)\prod_{\widetilde{Q}\in\mathcal{Q}_{h}}\widetilde{f}_{\widetilde{Q}}(m+\widetilde{Q}(n)),\end{split}$
where $\widetilde{f}_{\widetilde{Q}}$ are some $1$-bounded functions and $\mathcal{Q}_{h}$ is the collection (of size $\leq 2|\mathcal{Q}|$ and degree $\leq\deg\mathcal{Q}$) given by
@@
where $\widetilde{f}_{\widetilde{Q},a}=\widetilde{f}_{\widetilde{Q}}(\cdot+a)$.
Since by assumption $Q([N])\subset[-CN^{d},CN^{d}]$, basic linear algebra gives that the coefficients of $Q$ are $\ll_{C,d}1$ in modulus. Then, by the mean value theorem, for any $\widetilde{Q}\in\mathcal{Q}_{h}$ we have
$\max_{n\in[N]}|\widetilde{Q}(n)|\ll|\widetilde{Q}(1)|+N\max_{y\in[1,N]}|\widetilde{Q}^{\prime}(y)|\ll_{\deg\mathcal{Q},C}N^{d_{h}},$
so for all $n\in[N]$ we have $\widetilde{Q}(n)\in[-C^{\prime}N^{d_{h}},C^{\prime}N^{d_{h}}]$ for some $C^{\prime}\ll_{\deg\mathcal{Q},C}1$. Now, by the induction assumption (and the fact that the functions $\widetilde{f}_{Q,jN^{d}_{h}}$ may be assumed to be supported on $[-(C^{\prime}+1)N^{d_{h}},(C^{\prime}+1)N^{d_{h}}]$), we conclude that for some natural number $s^{\prime}\ll_{|\mathcal{Q}|,\deg{Q}}$ we have
@@
5. Quantitative uniformity of multiplicative functions
In this section, we give quantitative bounds for the uniformity norms of the Möbius function and other multiplicative functions that we need for the proofs of the main theorems.
5.1. The Möbius function
The only arithmetic property we need of the Möbius function is encapsulated in the following recent result of Leng [25], building on [26] and improving on [32].
Lemma 5.1 (Quantitative $U^{k}$-uniformity of $\mu$).
Let $k\in\mathbb{N}$, $A>0$ and $N\geq 3$. Then we have
$\|\mu\|_{U^{k}[N]}\ll_{A}(\log N)^{-A}.$
@@ 5.1. The Möbius function
This follows by combining Leng’s result [25, Theorems 6] with [32, Theorem 2.5] (where we can use Siegel’s bound $q_{\textnormal{Siegel}}\gg_{A}(\log N)^{A}$) and the triangle inequality for the Gowers norms. ∎
@@
We mention that the weaker bound $\|\mu\|_{u^{k}[N]}\ll_{A}(\log N)^{-A}$, which turns out to be all that we need for the proof of Theorem 1.2 in the case of distinct degree polynomials, is much simpler to prove. Indeed, it follows from the method to bilinear exponential sums developed by Vinogradov [33].